OpenStudy (anonymous):

Linear Approximation (.99)^25

OpenStudy (anonymous):

think of f(x)=x^25 f'(x)=25x^24 linear approximation f(25)= f(1)+ f'(1)(0.99-1)=1-25*0.01=0.75 which is fairly close to the actual value: 0.777821

OpenStudy (anonymous):

f(25) that is f(1)

OpenStudy (anonymous):

Andras has it. you could also use \[f(x)=(1-x)^{25}\] \[f'(x)=-25(1-x)^{24}\] find the equation of the line tangent to the curve and then let x = .1

OpenStudy (anonymous):

Thanks!!! You guys are awesome!! :D